OpenStudy (anonymous):

Find all solutions for cosθ = √2 - cosθ?

OpenStudy (anonymous):

Help would be awesome! If I could even just get the formula to use, it would be great help. I'm really lost. Thank you.

OpenStudy (solomonzelman):

cosθ = √2 - cosθ add cosθ to both sides 2cosθ=√2 divide both sides by 2 cosθ=√2/2 What angle is this ? 

OpenStudy (anonymous):

@SolomonZelman - That's what I had! Thank you! We are supposed to convert the answer into radians, however.. Let me see. Would you mind checking my work afterwards?

OpenStudy (solomonzelman):

If I am online, and if I am good enough at math, b/c I am not.

OpenStudy (solomonzelman):

I will try if I am on.

OpenStudy (anonymous):

@solomonZelman - cosθ = √2 - cosθ 2cosθ=√2 cosθ=√2/2 θ = cos-1(√2/2) θ = π/4 + 2πn

OpenStudy (solomonzelman):

After you got cosθ=√2/2 and after looking at my post you will probably get it.

OpenStudy (anonymous):

OR θ = 45º + 360ºn

OpenStudy (solomonzelman):

Yes, so... the answer is ?

OpenStudy (solomonzelman):

the theta = ?

OpenStudy (anonymous):

That is the answer.

OpenStudy (anonymous):

I just figured it out :)

OpenStudy (anonymous):

I checked it and it's right. We may not be thinking about the same subject material.

OpenStudy (solomonzelman):

but one thing answer, very very short. What is it ?

OpenStudy (anonymous):

That is the answer.. θ = 45º + 360ºn

OpenStudy (solomonzelman):

I would just say that it is 45º but either way.

OpenStudy (anonymous):

Oh. I know, but my teacher wants it this way. I didn't mean to come off as rude. :-) Thanks for your help nonetheless.